# RoHOI: Robustness Benchmark for Human-Object Interaction Detection

### v1

@@ C.4 Geometric and Scene (G&S) Distortions
Moire Pattern (MP): Interference patterns resulting from fine texture overlap with the camera’s pixel grid (e.g., fabric or screens). This artifact captures the challenges posed by imaging repetitive textures or high-frequency patterns. Severity levels adjust the density and contrast of the interference patterns. (Details shown in Fig. 8, row $1$.)
@@
Screen Crack Artifacts (SC): Visual obstructions or patterns caused by photographing through cracked or damaged surfaces. This type is included to reflect realistic scenarios where images are captured through broken screens or fractured glass. Higher severity levels introduce progressively denser and larger crack patterns, further obscuring the image. (Details shown in Fig. 8, row $2$.)
@@
Elastic Transform (ET): Non-linear distortions caused by perspective changes, material deformations, or uneven air density. This artifact represents challenges encountered in scenes with flexible or moving objects. Higher severity levels correspond to greater displacement magnitudes, leading to noticeable warping. (Details shown in Fig. 8, row $3$.)
@@
Perspective Distortion (PD): This artifact simulates geometric distortions from lens imperfections, such as barrel and pincushion distortion. Severity is increased by applying progressively stronger transformation matrices, exaggerating the curvature and altering the perceived structure of objects. (Details shown in Fig. 8, row $4$.)
@@
Pixelation (PIX): Blocky appearance due to low resolution or excessive digital zoom. This represents a key limitation in devices with constrained sensor capabilities or in images subjected to aggressive downscaling. Higher severity levels result in coarser pixel blocks, significantly reducing image detail. (Details shown in Fig. 8, row $5$.)
@@
Zoom Blur (ZB): Radial blur caused by rapid changes in focal length during image capture. Zoom blur is simulated by iteratively applying slight scaling transformations and averaging the results to create a progressive radial blur effect. Higher severity levels increase both the scaling intensity and the number of blended layers, leading to a stronger blur radiating outward from the center. This method ensures a smooth and natural-looking zoom effect without abrupt artifacts. (Details shown in Fig. 8, row $6$.)
C.5 Purpose of Inclusion

### v2

@@ Appendix C Corruption Types in the Dataset
Moire Pattern (MP): Interference patterns resulting from fine texture overlap with the camera’s pixel grid (e.g., fabric or screens). This artifact captures the challenges posed by imaging repetitive textures or high-frequency patterns. Severity levels adjust the density and contrast of the interference patterns. (Details shown in Figure 8, row $1$.)
@@
Screen Crack Artifacts (SC): Visual obstructions or patterns caused by photographing through cracked or damaged surfaces. This type is included to reflect realistic scenarios where images are captured through broken screens or fractured glass. Higher severity levels introduce progressively denser and larger crack patterns, further obscuring the image. (Details shown in Figure 8, row $2$.)
@@
Elastic Transform (ET): Non-linear distortions caused by perspective changes, material deformations, or uneven air density. This artifact represents challenges encountered in scenes with flexible or moving objects. Higher severity levels correspond to greater displacement magnitudes, leading to noticeable warping. (Details shown in Figure 8, row $3$.)
@@
Perspective Distortion (PD): This artifact simulates geometric distortions from lens imperfections, such as barrel and pincushion distortion. Severity is increased by applying progressively stronger transformation matrices, exaggerating the curvature and altering the perceived structure of objects. (Details shown in Figure 8, row $4$.)
@@
Pixelation (PIX): Blocky appearance due to low resolution or excessive digital zoom. This represents a key limitation in devices with constrained sensor capabilities or in images subjected to aggressive downscaling. Higher severity levels result in coarser pixel blocks, significantly reducing image detail. (Details shown in Figure 8, row $5$.)
@@
Zoom Blur (ZB): Radial blur caused by rapid changes in focal length during image capture. Zoom blur is simulated by iteratively applying slight scaling transformations and averaging the results to create a progressive radial blur effect. Higher severity levels increase both the scaling intensity and the number of blended layers, leading to a stronger blur radiating outward from the center. This method ensures a smooth and natural-looking zoom effect without abrupt artifacts. (Details shown in Figure 8, row $6$.)
Purpose of Inclusion